Market demand function is given by Qd = 180 – 2P

Market supply function is given by Qs = ‐ 15 + P

Determine producer surplus, consumer surplus and total economic surplus at the equilibrium price. Calculate the DWL that would result from a price floor imposed at a level of 72. 


1
Expert's answer
2021-03-22T08:07:35-0400

Consumer surplus is the difference between what a consumer is willing to pay and what they actually pay for a good or service.

On the other hand, producer surplus is the difference between the actual price of a good or service–the market price–and the lowest price a producer would be willing to accept for a good.

Consumer surplus

At equilibrium; quantity demanded equal the quantity supplied

"Qd=180-2P"

"Qs=-15+P"

"180-2P=-15+P"

"180+15=P+2P"

"P=65"

"Qd=180-2(65)"

"Qd=50"

At price 0 the quantity demanded is 180 and at quantity 0 units the Price is 90

At 0 units of quantity supplied the price is 15

Consumer Surplus (CS) = 1/2*b*h

Consumer surplus (CS) ="1\/2*50*(90-65)"

"CS=1\/2*50*25"

"CS=625"

Producer Surplus (PS) = "1\/2*b*h"

"PS=1\/2*50*50"

"PS =1,250"

Total Economic Surplus (TEC) = Consumer Surplus + Producer Surplus

"TEC=625+1,250"

"TEC=1,875"

Deadweight Loss (DWL) as result of price floor imposed at 72

"Qd=180-2(72)"

"Qd =36" implying that as price increases the quantity demanded decreases.

At "Qd=36" Then the price for the supply will reduce to 51

"36=-15+P"

"P=51"

Dead Weight Loss (DWL) = 1/2* Price difference*Quantity Difference

"DWL=1\/2*(P2-P1)*(Qe-Q2)"

"DWL=1\/2*(72-51)*(50-36)"

"DWL=1\/2*21*14"

"DWL=147"